OpenStudy (anonymous):

PLEASE HELP!! What is the slope-intercept form of the function that contains the points (3, 1) and (2, 9)? Y=__x+__

OpenStudy (anonymous):

\[slope = \frac{ y2-y1 }{ x2-x1}\]

OpenStudy (anonymous):

SO the slope is -8

OpenStudy (anonymous):

y=mx+b slope = m

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I know that I am just confused on how to get B

OpenStudy (anonymous):

Ok...use the following\[(y-y1)=8(x-x1)\]

OpenStudy (anonymous):

Let y1 be either one of the points on the line (3,1) or (2,9)

OpenStudy (anonymous):

okay...I will try

OpenStudy (anonymous):

Is it -24

OpenStudy (anonymous):

Your answer needs to be something that looks like y=mx+b Use one of the ordered pairs and put them in this equation for x and y. Use the slope you found above for m. Now solve the equation for b.

OpenStudy (anonymous):

Okay so the answer is y=-8+-24

OpenStudy (anonymous):

\[(y-1) = -8(x-3)\]

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

You still need to get the -1 off the left side of the equation

OpenStudy (anonymous):

So add 1 to both sides and what do yu have?

OpenStudy (anonymous):

I am lost i am so sorry

OpenStudy (anonymous):

Sorry should be y-1= -8x +24 now add 1 to both sides

OpenStudy (anonymous):

Ok we know the slope is -8 correct.

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

The formula for the line is (y-y1) = m(x-x1) with m being the slope.

OpenStudy (anonymous):

So we have (y-y1) = -8 ( x-x1)

OpenStudy (anonymous):

You see this?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

OK now we have (y-y1)= -8 (x-x1). You now plug in the values from the points on the line for x1 and y1. I will use the point (3,1). So it will look like this (y-1) = -8(x-3) Still with me?

OpenStudy (anonymous):

Yes i understand

OpenStudy (anonymous):

Great now you need to distribute the right side of the equation to get (you did this earlier) now we have y-1 = -8x+24 Still OK with everything?

OpenStudy (anonymous):

Yeah I am okay

OpenStudy (anonymous):

The formula is y=mx+b....we have y-1 = -8x+24..so we have to get rid of the -1 on the left side of the equation. To do that you will add 1 to both sides of the equation and what do you end up with?

OpenStudy (anonymous):

You end up with y=-9x+24

OpenStudy (anonymous):

No you cannot add 1 to 8x you add it to 24

OpenStudy (anonymous):

okay sorry

OpenStudy (anonymous):

So it would be y=-8+-25

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then you divide both sides right

OpenStudy (anonymous):

No the answer is y=-8x+25

OpenStudy (anonymous):

okay THANK YOU SO MUCH

OpenStudy (anonymous):

YW